# ZPressor: Bottleneck-Aware Compression for Scalable Feed-Forward 3DGS

### v2

@@ 4.1 Experimental Settings
Implementation details. We use the same computing resources to train the baseline and our method. Due to the memory limit, we use 6 context views for DepthSplat and MVSplat, and 4 context views for pixelSplat. For all of our experiments, we adopted the same learning rate as the baseline, utilized the AdamW optimizer, and trained the models for 100,000 steps on A800 GPUs. Following the setting of the baseline, we use the $256\times 256$ input resolution on RE10K, and $256\times 448$ input resolution on DL3DV. All training losses match those of the baseline, with no additional data or regularization introduced. More details are provided in Appendix B.
4.2 SoTA Comparisons and IB Analysis
@@
Comparisons with SoTA models. We train all models on DL3DV and RE10K using 12 input views with 6 anchor views set to our ZPressor, and evaluate them under varying numbers of input views ranging from 8 to 36. As shown in Tab. 1 and Tab. 2, integrating ZPressor into DepthSplat, MVSplat, and pixelSplat consistently improves their performance across all input view settings and evaluation metrics, demonstrating the effectiveness of our approach.
Notably, the performance gain becomes more significant as the number of input views increases. This is because existing feed-forward 3DGS models struggle with dense inputs due to information overload, leading to performance degradation. In contrast, ZPressor mitigates this issue by compressing the input through redundancy suppression while preserving essential information, improving model robustness, and maintaining strong performance under dense input settings. Moreover, we observe that pixelSplat fails to run with more than 8 input views due to out-of-memory (OOM) caused by the large number of predicted pixel-aligned 3D Gaussians. In contrast, our ZPressor helps merge input information, reducing the number of predicted Gaussians and enabling testing with up to 36 views. These observations are further supported by the qualitative comparisons shown in Fig. 3 and Fig. 4, where DepthSplat and MVSplat exhibit noticeable artifacts under 36 input views, whereas their ZPressor-augmented versions produce significantly cleaner renderings. More results are provided in Appendix D.
[12]
@@
Analysis of model efficiency. Compressing the input view information not only improves robustness and performance, but also enhances efficiency. To validate this, we compare the model efficiency of DepthSplat with and without ZPressor under 480P resolution, evaluating the number of 3D Gaussians, the test-time inference latency, and the peak memory usage. As shown in Fig. 5, the benefits of integrating ZPressor are clear. In particular, as the number of context views increases, the baseline model’s predicted 3D Gaussian numbers, memory usage, and inference time all grow linearly, whereas ZPressor helps maintain stable resource consumption across all aspects.
@@
We can then analysis the effect of bottleneck constraint (the number of anchor views) under scenes with varying information content by adjusting the frame distance between input views.
As shown in Sec. 4.2, we conduct experiments under two settings, Context Gap 50 (CG50, in blue) and Context Gap 100 (CG100, in pink), where context gap refers to the frame distance between input views. CG100 thus contains more information than CG50. Under CG50, increasing the number of anchor views from 7 to 9 leads to a performance drop, suggesting that 7 clusters are sufficient and additional ones introduce redundancy. In contrast, under CG100, increasing anchors from 7 to 9 improves performance,indicating a higher information bottleneck for scenes with more information. These results highlight the effectiveness of our ZPressor in implementing the IB principle and show that the information bottleneck is critical in balancing compression and information preservation.
Analysis of information fusion. To further confirm that our ZPressor effectively fuses information from support views into anchor views rather than simply discarding it, we conduct experiments by varying the fusion strategy. As shown in Sec. 4.2, removing the information fusion step (“w/o fusion”) leads to a performance drop, highlighting the importance of fusing support views into anchor views. To ensure that the introduction of support information does result in a performance improvement, we conduct a control experiment by fusing repeated anchor views instead of support views (“fuse anchors”). Since this does not introduce new information, its performance is lower than our default setting, which achieves the best results by fusing complementary information. These comparisons further validate that our design effectively implements the IB principle.
4.3 Ablation Study
As mentioned at the end of Sec. 3, the default ZPressor uses several stacked attention blocks, each combining self-attention and cross-attention. This section presents an ablation study to validate our design.
We report results on DL3DV with 12 input views, following the setting in Tab. 1, using DepthSplat as the backbone. As shown in Tab. 3, removing the stacking design and using only one block (“w/o multi-blocks”) slightly degrades performance, suggesting that stacking improves fusion of information from support to anchor views. Additionally, removing the self-attention (“w/o self attention”) also reduce performance, showing that self-attention complements cross-attention by enhancing internal feature interactions. Overall, all variants of ZPressor outperform the DepthSplat baseline, confirming the existence of an information bottleneck in feed-forward 3DGS models and the effectiveness of our ZPressor in addressing it.
@@ 5 Conclusion
We have provided a fundamental analysis of the model capacity limitations in existing feed-forward 3DGS models through the lens of the Information Bottleneck principle. Building on this insight, we introduced ZPressor, a lightweight, architecture-agnostic module that efficiently compresses multi-view inputs, enabling models to overcome inherent limitations and scale to handle more input views. We validated our ZPressor by integrating it into several representative feed-forward 3DGS models. Our experiments on several large-scale benchmarks demonstrate that ZPressor not only consistently improves the performance of existing models under moderate view settings, but also helps them maintain competitive efficiency under denser inputs. We believe that ZPressor significantly enhances the scalability and practicality of feed-forward 3DGS models, opening the door to more effective applications in real-world scenarios.
@@ Appendix A More Experimental Analysis
Cross dataset generalization. Following MVSplat [8], we conducted experiments using a pretrained model on the RealEstate10K (RE10K) dataset [21] (as detailed in Tab. 2) and tested its performance on the ACID dataset [67] to evaluate the generalization capabilities of our proposed ZPressor across diverse datasets. As demonstrated in Table A, MVSplat with ZPressor exhibits remarkable efficacy in cross-dataset generalization. Notably, this performance advantage becomes progressively more pronounced with an increasing number of input views.
@@ Appendix B More Implementation Details
Network architectures. In Algorithm 1, we provide a detailed description of how ZPressor is integrated into existing feed-forward 3D Gaussian Splatting (3DGS) frameworks [7, 12, 8]. Initially, we select anchor views and their corresponding support views following Algorithm 2 and Eq. (5). The features associated with these views are then processed by an attention-based network. This network is composed of 6 structurally identical blocks, wherein each block encompasses a cross-attention layer, a self-attention layer, and an MLP layer. The cross-attention mechanism operates by employing the anchor features as query, while the support features provide the key and value. Subsequent to this fusion, the resulting features are further refined by the self-attention and MLP layers.
Algorithm 1 Overview of Feed-Forward 3DGS framework with ZPressor
@@
     $\mathcal{Z}\leftarrow\text{Cross-Attn}(Q,K,V)$, where $Q\leftarrow\mathcal{Z}\quad K,V\leftarrow\mathcal{X}_{\mathrm{support}}$
     $\mathcal{Z}\leftarrow\text{Self-Attn}(Q,K,V)$, where $Q,K,V\leftarrow\mathcal{Z}$
@@
Set of view camera positions $\mathcal{T}=\{\mathbf{T}_{1},\mathbf{T}_{2},...,\mathbf{T}_{K}\}$, Number of anchor views $N$
@@ Appendix C Limitation and Societal Impacts
Limitation analysis. As discussed in Sec. 5, ZPressor exhibits limitations when processing scenarios with an extremely high density of input views. Specifically, its efficacy in compressing the information from such dense views through a limited set of anchor views is diminished. To illustrate this, we conducted an experiment on DepthSplat [12] integrated with ZPressor, using approximately 500 images as input. As depicted in Fig. A, the quality of the rendered novel views was perceptibly affected, which can be attributed to an insufficient number of Gaussian primitives to adequately represent the scene under these dense input conditions.
@@ Appendix D More Visual Comparisons
This section provides additional qualitative comparison results. We present further visualizations for DepthSplat [12] on the DL3DV [17] and MVSplat [8] on the RE10K [21] in Fig. B and Fig. F, with our ZPressor.
Furthermore, to illustrate how ZPressor performs with dense input views, we showcase comparative results. For DepthSplat [12], comparisons between the original framework and DepthSplat augmented with ZPressor are presented for scenarios with 24, 16, and 12 input views in Fig. C, Fig. D, and Fig. E. Similarly, for MVSplat [8], visual comparisons between the original framework and MVSplat integrated with ZPressor are displayed for inputs of 24, 16, and 8 views in Fig. G, Fig. H, and Fig. I. The corresponding quantitative results for these multi-view experiments can be found in Tab. 2.
[12]

### v4

@@ 4.1 Experimental Settings
Implementation details. We use the same computing resources to train the baseline and our method. Due to the memory limit, we use 6 context views for DepthSplat and MVSplat, and 4 context views for pixelSplat. For all of our experiments, we adopted the same learning rate as the baseline, utilized the AdamW optimizer, and trained the models for 100,000 steps on A800 GPUs. Following the setting of the baseline, we use the $256\times 256$ input resolution on RE10K, and $256\times 448$ input resolution on DL3DV. All training losses match those of the baseline, with no additional data or regularization introduced.
4.2 SoTA Comparisons and IB Analysis
Comparisons with SoTA models. We train all models on DL3DV and RE10K using 12 input views with 6 anchor views set to our ZPressor, and evaluate them under varying numbers of input views ranging from 8 to 36. As shown in Tab.˜1 and Tab.˜2, integrating ZPressor into DepthSplat, MVSplat, and pixelSplat consistently improves their performance across all input view settings and evaluation metrics, demonstrating the effectiveness of our approach.
Notably, the performance gain becomes more significant as the number of input views increases. This is because existing feed-forward 3DGS models struggle with dense inputs due to representation overload, leading to performance degradation. In contrast, ZPressor mitigates this issue by compressing the input through redundancy suppression while preserving essential information, improving model robustness, and maintaining strong performance under dense input settings. Moreover, we observe that pixelSplat fails to run with more than 8 input views due to out-of-memory (OOM) caused by the large number of predicted pixel-aligned 3D Gaussians. In contrast, our ZPressor helps merge input information, reducing the number of predicted Gaussians and enabling testing with up to 36 views. These observations are further supported by the qualitative comparisons shown in Fig.˜3 and Fig.˜4, where DepthSplat and MVSplat exhibit noticeable artifacts under 36 input views, whereas their ZPressor-augmented versions produce significantly cleaner renderings.
Refer to caption[12]
@@ 4.2 SoTA Comparisons and IB Analysis
tableAnalysis of Information Fusion. “default” denotes our setting where support views are fused into anchor views. “w/o fusion” removes the fusion step, and “fuse anchors” fuses repeated anchor views instead. “default” performs best, indicating that ZPressor improves performance by effectively fusing complementary information from the support views.
@@
default 24.30 0.821 0.146
fuse anchors 24.23 0.817 0.148
w/o fusion 23.80 0.810 0.162
Analysis of model efficiency. Compressing the input view information not only improves robustness and performance, but also enhances efficiency. To validate this, we compare the model efficiency of DepthSplat with and without ZPressor under 480P resolution, evaluating the number of 3D Gaussians, the test-time inference latency, and the peak memory usage. As shown in Fig.˜5, the benefits of integrating ZPressor are clear. In particular, as the number of context views increases, the baseline model’s predicted 3D Gaussian numbers, memory usage, and inference time all grow linearly, whereas ZPressor helps maintain stable resource consumption across all aspects.
@@
We can then analysis the effect of bottleneck constraint (the number of anchor views) under scenes with varying information content by adjusting the frame distance between input views. As shown in Sec.˜4.2, we conduct experiments under two settings, Context Gap 50 (CG50, in blue) and Context Gap 100 (CG100, in pink), where context gap refers to the frame distance between input views. When the scene’s information content is relatively low (e.g., small camera baseline, similar views, as proxied by a small CG like 50), a smaller number of anchor views (e.g., 7) is already sufficient to capture the essential scene information. Adding more anchor views in such a scenario might introduce redundancy or ambiguity, as these additional anchors might not observe genuinely new regions but rather re-observe already covered areas from slightly different perspectives. This could lead to a less compact and potentially noisier latent representation, hence the performance drop. In contrast, for scenes with higher information content (larger CG like 100), more anchor views are beneficial as they help cover more diverse perspectives and capture richer scene details. These results highlight the effectiveness of our ZPressor as an instantiation of the IB principle and show that the information bottleneck is critical in balancing compression and information preservation.
@@
Analysis of information fusion. To further confirm that our ZPressor effectively fuses information from support views into anchor views rather than simply discarding it, we conduct experiments by varying the fusion strategy. As shown in Sec.˜4.2, removing the information fusion step (“w/o fusion”) leads to a performance drop, highlighting the importance of fusing support views into anchor views. To ensure that the introduction of support information does result in a performance improvement, we conduct a control experiment by fusing repeated anchor views instead of support views (“fuse anchors”). Since this does not introduce new information, its performance is lower than our default setting, which achieves the best results by fusing complementary information. These comparisons further validate that our design effectively implements the IB principle.
4.3 Ablation Study
As mentioned at the end of Sec.˜3, the default ZPressor uses several stacked attention blocks, each combining self-attention and cross-attention. This section presents an ablation study to validate our design.
We report results on DL3DV with 12 input views, following the setting in Tab.˜1, using DepthSplat as the backbone. As shown in Tab.˜3, removing the stacking design and using only one block (“w/o multi-blocks”) slightly degrades performance, suggesting that stacking improves fusion of information from support to anchor views. Additionally, removing the self-attention (“w/o self attention”) also reduce performance, showing that self-attention complements cross-attention by enhancing internal feature interactions. Overall, all variants of ZPressor outperform the DepthSplat baseline, confirming the existence of an information bottleneck in feed-forward 3DGS models and the effectiveness of our ZPressor in addressing it.
@@ 5 Conclusion
We have provided a fundamental analysis of the model capacity limitations in existing feed-forward 3DGS models through the lens of the Information Bottleneck principle. Building on this insight, we have introduced ZPressor, a lightweight, architecture-agnostic module that efficiently compresses multi-view inputs, enabling models to overcome inherent limitations and scale to handle more input views. We have validated our ZPressor by integrating it into several representative feed-forward 3DGS models. Our experiments on several large-scale benchmarks have demonstrated that ZPressor not only consistently improves the performance of existing models under moderate view settings, but also helps them maintain competitive efficiency under denser inputs. We believe that ZPressor significantly enhances the scalability and practicality of feed-forward 3DGS models, opening the door to more effective applications in real-world scenarios.
@@
Acknowledgements. We thank Biao Wu for his enlightening discussions and inspiration on network architecture.
@@ Appendix A More Experimental Analysis
A.1 Cross Dataset Generalization
Following MVSplat [8], we conducted experiments using a pretrained model on the RealEstate10K (RE10K) dataset [21] (as detailed in Tab.˜2) and tested its performance on the ACID dataset [68] to evaluate the generalization capabilities of our proposed ZPressor across diverse datasets. As demonstrated in Table A, MVSplat with ZPressor exhibits remarkable efficacy in cross-dataset generalization. Notably, this performance advantage becomes progressively more pronounced with an increasing number of input views.
@@ A.1 Cross Dataset Generalization
A.2 Other Selection Strategies
In addition to our final choice of FPS-based selection strategy (Fig.˜2 Step 1), we also experimented with other anchor views selection strategies. Below are detailed implementation descriptions for each strategy and a comparison of all strategies’ performance in Tab.˜B.
Overlap-based selection. We try overlap-based selection by computing a pairwise overlap matrix between camera views. The overlap is obtained by projecting dense pixel rays from one view into another and averaging the proportion of pixels that fall within the target image plane. The final overlap score between two views is the minimum of the bidirectional projections. We then apply a greedy vectorized algorithm that iteratively selects views with the lowest average overlap to the already chosen set, ensuring maximal scene coverage with minimal redundancy. This procedure produces both the anchor indices and the cluster assignments for the remaining views.
Pose-free settings. In the absence of camera poses, we cluster learned per-view tokens instead of 3D positions. We use K-Means in the feature space of global view embeddings. Each token is assigned to its nearest cluster center, and the closest token to the mean of each cluster is selected as the anchor. This pose-free clustering provides a practical alternative for unposed datasets, as the anchor set is derived entirely from image content rather than camera geometry.
Table B: Quantitative comparison of different selection strategies integrated with DepthSplat [12]. All variants are evaluated with 36 test views on DL3DV [17].
[…]
DepthSplat +0.01 19.23+0.01 +0.01 0.666+0.01 +0.01 0.286+0.01
DepthSplat + Ours (overlap-based) +0.01 21.49 +2.26 +0.001 0.727 +0.061 +0.001 0.194 -0.092
DepthSplat + Ours (pose-free) +0.01 22.81 +3.58 +0.001 0.791 +0.125 +0.001 0.174 -0.112
DepthSplat + Ours (K-Means-based) +0.01 22.84 +3.61 +0.001 0.789 +0.123 +0.001 0.175 -0.111
DepthSplat + Ours +0.01 23.88 +4.65 +0.001 0.815 +0.149 +0.001 0.150 -0.136
K-Means-based selection. K-Means-based selection groups view positions directly. Given the 3D camera centers of all input views, we apply K-Means with a fixed number of groups and identify the anchor of each cluster as the view closest to its centroid. Each remaining view is assigned to its nearest centroid, producing compact and spatially coherent groups. This procedure is fully deterministic given the random seed and provides stable anchors that reflect the geometric distribution of input cameras.
A.3 Comparison with Confidence-based Pruning
We compare ZPressor against confidence-based pruning after predicting Gaussians in MVSplat [8]. Confidence pruning removes input views based on their predicted reliability scores, with a fixed pruning ratio controlling the proportion of views retained. When maintaining the same number of Gaussians, the aggressive removal of views leads to a substantial performance drop due to insufficient multi-view support. At a pruning ratio of $0.5$, the method achieve the best accuracy, but the gain remains limited because the retained subset does not guarantee spatial coverage of the scene. In contrast, ZPressor consistently compresses all input views into compact latent anchors while maintaining balanced coverage, which yields superior accuracy without discarding information.
Table C: Comparison with confidence-based pruning (CP) in MVSplat [8] under 24 views on RE10K. For a fair comparison, the prune ratio for CP is set to $0.75$ to yield a same number of Gaussians as ZPressor, and we also manually adjust the prune ratio to $0.5$ to achieve the best performance. All models are re-trained.
[…]
MVSplat +0.01 25.00+0.01 +0.001 0.871+0.001 +0.001 0.137+0.001
MVSplat + CP (prune_ratio=0.75prune\_ratio=0.75) +0.01 21.15 -3.85 +0.001 0.816 -0.055 +0.001 0.190 +0.053
MVSplat + CP (prune_ratio=0.5prune\_ratio=0.5) +0.01 26.94 +1.94 +0.001 0.886 +0.015 +0.001 0.130 -0.007
MVSplat + Ours +0.01 27.49 +2.49 +0.001 0.895 +0.024 +0.001 0.111 -0.026
A.4 Ablation Study of IB-Loss
We conduct an ablation study to evaluate the role of the compression term derived from the Information Bottleneck formulation. When the compression term is removed ($\beta=0$), ZPressor still improves significantly over the baseline DepthSplat, confirming that the architectural design alone provides strong benefits. Introducing a small but non-zero coefficient ($\beta=10^{-5}$) further encourages compact latent representations and yields the best balance between distortion and fidelity. This demonstrates that the IB-inspired loss can serve as a lightweight regularizer to assist compression module learning in achieving better compression results.
Table D: Ablation of the IB-inspired compression loss on DepthSplat [12] with 36 views on DL3DV [17]. The compression term provides additional regularization that improves overall performance.
[…]
DepthSplat +0.01 19.23+0.01 +0.001 0.666+0.001 +0.001 0.286+0.001
DepthSplat + Ours ($\beta=0$) +0.01 23.43 +4.20 +0.001 0.806 +0.140 +0.001 0.165 -0.121
DepthSplat + Ours ($\beta=10^{-5}$) +0.01 23.88 +4.65 +0.001 0.815 +0.149 +0.001 0.150 -0.136
@@ Appendix B More Implementation Details
Network architectures. In Algorithm˜1, we provide a detailed description of how ZPressor is integrated into existing feed-forward 3D Gaussian Splatting (3DGS) frameworks [7, 12, 8]. Initially, we select anchor views and their corresponding support views following Algorithm˜2 and Eq. (5). The features associated with these views are then processed by an attention-based network. This network is composed of 6 structurally identical blocks, wherein each block encompasses a cross-attention layer, a self-attention layer, and an MLP layer. The cross-attention mechanism operates by employing the anchor features as query, while the support features provide the key and value. Subsequent to this fusion, the resulting features are further refined by the self-attention and MLP layers.
Algorithm 1 Overview of Feed-Forward 3DGS framework with ZPressor.
@@
  $\mathcal{Z}\leftarrow\text{Attention}(Q,K,V)$, where $Q\leftarrow\mathcal{Z}\quad K,V\leftarrow\mathcal{X}_{\mathrm{support}}$
  $\mathcal{Z}\leftarrow\text{Attention}(Q,K,V)$, where $Q,K,V\leftarrow\mathcal{Z}$
@@
Set of view camera positions $\mathcal{T}=\{\mathbf{T}_{1},\mathbf{T}_{2},...,\mathbf{T}_{K}\}$, Number of anchor views NN
@@ Appendix C Limitation and Societal Impacts
Limitation analysis. As discussed in Sec.˜5, ZPressor exhibits limitations when processing scenarios with an extremely high density of input views. Specifically, its efficacy in compressing the information from such dense views through a limited set of anchor views is diminished. To illustrate this, we conducted an experiment on DepthSplat [12] integrated with ZPressor, using approximately 500 images as input. As depicted in Fig.˜A, the quality of the rendered novel views was perceptibly affected, which can be attributed to an insufficient number of Gaussian primitives to adequately represent the scene under these dense input conditions.
@@ Appendix D More Visual Comparisons
This section provides additional qualitative comparison results. We present further visualizations for DepthSplat [12] on the DL3DV [17] and MVSplat [8] on the RE10K [21] in Fig.˜B and Fig.˜F, with our ZPressor.
Furthermore, to illustrate how ZPressor performs with dense input views, we showcase comparative results. For DepthSplat [12], comparisons between the original framework and DepthSplat augmented with ZPressor are presented for scenarios with 24, 16, and 12 input views in Fig.˜C, Fig.˜D, and Fig.˜E. Similarly, for MVSplat [8], visual comparisons between the original framework and MVSplat integrated with ZPressor are displayed for inputs of 24, 16, and 8 views in Fig.˜G, Fig.˜H, and Fig.˜I. The corresponding quantitative results for these multi-view experiments can be found in Tab.˜2.
Refer to caption[12]